# A unified convention for achievement positional games

### v1

@@ 5 Conclusion
We have introduced achievement positional games, a new convention for positional games where the players try to fill different edges. We have established some of their general properties, which do not significantly differ from those of Maker-Maker games, and obtained complexity results for almost all edge sizes. A corollary is that the Maker-Maker convention is PSPACE-complete for positions that can be obtained from a hypergraph of rank 4 after just one round of play, which is the first known complexity result on this convention for edges of size between 3 and 5.
We have not determined the exact complexity of the cases $(p,q)\in\{(3,2),(4,2)\}$, even though we know they are NP-hard. The commonplace intuition within the community is that MakerBreaker(4) is PSPACE-complete, which would imply that AchievementPos($4$,$2$) also is. Conversely, AchievementPos($4$,$2$) could be settled first, as a way to approach the elusive hardness of MakerBreaker(4) by giving Breaker some edges of her own. The case of AchievementPos($3$,$2$) is different, as we already know that there is a complexity gap compared to MakerBreaker(3). A proof of either membership in NP or PSPACE-hardness would be compelling.
@@
As mentioned in the introduction, one can define the Maker-Breaker convention as the special case of achievement positional games where each of $(V,E_{L})$ and $(V,E_{R})$ is the transversal hypergraph of the other. This inclusion of the transversals in the input game is mathematically redundant, but it potentially changes everything from an algorithmic standpoint, as even hypergraphs with bounded rank can have an exponential number of minimal transversals. Do Maker-Breaker games become easy if both players’ winning sets are part of the input rather than solely Maker’s? Answering this question would help understanding what makes the Maker-Breaker convention hard (which it is for 5-uniform hypergraphs [Koe25]): does it come from the hardness of computing the transversal hypergraph, or would it still be hard if we were given all the transversals.
@@
Like all positional games, achievement positional games can be generalized to CNF formulas in place of hypergraphs. We would have two formulas, one for Left and one for Right, on the same set of variables. Left and Right takes turns choosing a variable and a valuation (T or F) for that variable. Whoever first falsifies their own formula wins. Achievement positional games correspond to the special case where Left’s formula is positive and Right’s formula is negative, so that Left (resp. Right) always chooses the valuation F (resp. T). As AchievementPos($2$,$2$) is in P, it would be interesting to see if the case where all clauses have size at most 2 is tractable.
@@
Another natural prospect would be to define avoidance positional games, where whoever first fills an edge of their color loses. Since the Avoider-Avoider convention is already PSPACE-complete for 2-uniform hypergraphs [FGM+15], the complexity aspects would not be as interesting. However, an analogous study to that of Section 3 could be performed to better understand general properties of avoidance games.
@@ Appendix A Examples for the outcome of a disjoint union
Some of them use the following construction, which can be seen as a generalization of the butterfly for any number of moves needed to win. We define the achievement positional game $\mathcal{W}_{k}^{L}=(V,E_{L},E_{R})$ where $V=\{u,v_{1},\ldots,v_{2k-2}\}$, $E_{L}=\{S\cup\{u\}\mid S\subseteq\{v_{1},\ldots,v_{2k-2}\},|S|=k-1\}$ and $E_{R}=\varnothing$. The idea is that $\mathcal{W}_{k}^{L}$ is a game with outcome $\mathcal{L}^{-}$, on which Left needs $k$ moves to win as first player (by picking $u$ then picking arbitrary vertices), and which Right can entirely destroy as first player by picking $u$ herself. Similarly, we define $\mathcal{W}_{k}^{R}$ by swapping $E_{L}$ with $E_{R}$.
@@
  • We have $o(\mathcal{G})=\mathcal{N}$. Indeed, Left wins in two moves as first player thanks to $\mathcal{W}_{2}^{L}$, and Right wins as first player by destroying $\mathcal{W}_{2}^{L}$ and then winning thanks to whichever of $\mathcal{W}_{5}^{R}$ or $\mathcal{W}_{7}^{R}$ is left intact.
  • We have $o(\mathcal{G}^{\prime})=\mathcal{N}$. Indeed, let $u$ be the universal vertex of the “T”-shaped part. Left wins in two moves as first player by picking $u$, and Right wins as first player by picking $u$ (which forces Left’s answer) and then winning thanks to $\mathcal{W}_{3}^{R}$.
  • We have $o(\mathcal{G}\cup\mathcal{G}^{\prime})=\mathcal{L}^{-}$. Indeed, Left wins in two moves as first player by picking $u$ for example. Now, suppose that Right starts. As she cannot ensure a win in less than three moves, she must start by destroying both of Left’s threats of winning in two moves. For this, she must pick $u$ first (which forces Left’s answer) and then destroy $\mathcal{W}_{2}^{L}$. After that, it is Left’s turn, but he cannot win in less than four moves while Right is threatening to win in three moves. Therefore, Left must destroy $\mathcal{W}_{3}^{R}$. According to the same reasoning, the game continues with Right destroying $\mathcal{W}_{4}^{L}$, Left destroying $\mathcal{W}_{5}^{R}$, Right destroying $\mathcal{W}_{6}^{L}$ and Left destroying $\mathcal{W}_{7}^{R}$. In conclusion, the game ends in a draw.

### v2

@@ 5 Conclusion
We have introduced achievement positional games, a new convention for positional games where the players try to fill different edges. We have established some of their general properties, which do not significantly differ from those of Maker-Maker games, and obtained complexity results for all edge sizes which were not already settled by previous results on positional games.
One of the motivations behind achievement positional games is to have a framework for intermediate positions of the Maker-Maker convention, since edges in which a player has picked a vertex can only be filled by that same player from there on. This idea has already materialized, in two ways. First of all, since the publication of the short conference version of this work [9], our approach has been used to show PSPACE-completeness of (starting positions of) the Maker-Maker convention on hypergraphs of rank 4 [10]. This leaves hypergraphs of rank 3 as the only open case, and our Corollary 4.8 is the first step made towards solving that case, as it states that the Maker-Maker convention is PSPACE-complete for positions that can be obtained from a 3-uniform hypergraph after just one round of play.
@@
As mentioned in the introduction, one can define the Maker-Breaker convention as the special case of achievement positional games where each of $(V,E_{L})$ and $(V,E_{R})$ is the transversal hypergraph of the other. This inclusion of the transversals in the input game is mathematically redundant, but it is potentially significant from an algorithmic standpoint, as even hypergraphs with bounded rank can have an exponential number of minimal transversals. Do Maker-Breaker games become easy if both players’ winning sets are part of the input rather than solely Maker’s? Answering this question would help understanding what makes the Maker-Breaker convention hard (which it is for 4-uniform hypergraphs [11]): does it come from the hardness of computing the transversal hypergraph, or would it still be hard if we were given all the transversals.
@@
Like all positional games, achievement positional games can be generalized to CNF formulas in place of hypergraphs. We would have two formulas, one for Left and one for Right, on the same set of variables. Left and Right takes turns choosing a variable and a valuation (T or F) for that variable. Whoever first falsifies their own formula wins. Achievement positional games correspond to the special case where Left’s formula is positive and Right’s formula is negative, so that Left (resp. Right) always chooses the valuation F (resp. T). As AchievementPos(22,22) is in P, it would be interesting to see if the case where all clauses have size at most 2 is tractable.
@@
Another natural prospect would be to define avoidance positional games, where whoever first fills an edge of their color loses. Since the Avoider-Avoider convention is already PSPACE-complete for 2-uniform hypergraphs [7], the complexity aspects would not be as interesting. However, an analogous study to that of Section 3 could be performed to better understand general properties of avoidance games.
@@ Appendix A Examples for the outcome of a disjoint union
Some of them use the following construction, which can be seen as a generalization of the butterfly for any number of moves needed to win. We define the achievement positional game $\mathcal{W}_{k}^{L}=(V,E_{L},E_{R})$ where $V=\{u,v_{1},\ldots,v_{2k-2}\}$, $E_{L}=\{S\cup\{u\}\mid S\subseteq\{v_{1},\ldots,v_{2k-2}\},|S|=k-1\}$ and $E_{R}=\varnothing$. The idea is that $\mathcal{W}_{k}^{L}$ is a game with outcome $\mathcal{L}^{-}$, on which Left needs $k$ moves to win as the first player (by picking $u$ then picking arbitrary vertices), and which Right can entirely destroy as the first player by picking $u$ herself. Similarly, we define $\mathcal{W}_{k}^{R}$ by swapping $E_{L}$ with $E_{R}$.
@@
  • We have $o(\mathcal{G})=\mathcal{N}$. Indeed, Left wins in two moves as the first player thanks to $\mathcal{W}_{2}^{L}$, and Right wins as the first player by destroying $\mathcal{W}_{2}^{L}$ and then winning thanks to whichever of $\mathcal{W}_{5}^{R}$ or $\mathcal{W}_{7}^{R}$ is left intact.
  • We have $o(\mathcal{G}^{\prime})=\mathcal{N}$. Indeed, let $u$ be the vertex of degree 3 in the “T”-shaped part. Left wins in two moves as the first player by picking $u$, and Right wins as the first player by picking $u$ (which forces Left’s answer) and then winning thanks to $\mathcal{W}_{3}^{R}$.
  • We have $o(\mathcal{G}\cup\mathcal{G}^{\prime})=\mathcal{L}^{-}$. Indeed, Left wins in two moves as the first player by picking $u$ for example. Now, suppose that Right starts. As she cannot ensure a win in less than three moves, she must start by destroying both of Left’s threats of winning in two moves. For this, she must pick $u$ first (which forces Left’s answer) and then destroy $\mathcal{W}_{2}^{L}$. After that, it is Left’s turn, but he cannot win in less than four moves while Right is threatening to win in three moves. Therefore, Left must destroy $\mathcal{W}_{3}^{R}$. According to the same reasoning, the game continues with Right destroying $\mathcal{W}_{4}^{L}$, Left destroying $\mathcal{W}_{5}^{R}$, Right destroying $\mathcal{W}_{6}^{L}$ and Left destroying $\mathcal{W}_{7}^{R}$. In conclusion, the game ends in a draw.